OpenStudy (anonymous):

Solve compound inequality. Write solution in interval notation. x+9<0 or 2x> -12 (these kill me)

OpenStudy (y2o2):

it has no solution

OpenStudy (y2o2):

x cant be grater than -6 and smaller than -9

OpenStudy (ash2326):

Jmallis it's easy, if we proceed step by step Let's solve the first equation x+9<0 subtract 9 from both sides x<-9 so x is \((-\infty,-9)\) now from the second equation 2x>-12 or x>-6 so from this equation x is \((-6, \infty)\) now it's given x+9<0 or x>-12 so we have to take union of the two ie x belongs to \((-\infty,-9)\) U \((-6, \infty)\)

OpenStudy (anonymous):

Thank-you, you think you can help me with another one ash2326

OpenStudy (ash2326):

yeah jmallis, post your question :)

OpenStudy (ash2326):

Jmallis , post your question as a new thread, that way you can get more people to respond:)

OpenStudy (anonymous):

\[-6\le2x-12\]\[\le18\]

OpenStudy (anonymous):

should be all together

OpenStudy (ash2326):

we have -6<=2x-12<=18 add 12 to all sides 6<=2x<=30 divide be 2 the whole inequation 3<=x<=10 As we can see x belongs to [3, 10] do notice that it's closed interval

OpenStudy (anonymous):

says to solve compund inequality written in compact form, get x alone in the middle part. Since a compound inequality is really two inequalities in one statement, perform same operations on all 3 parts of inequality

OpenStudy (ash2326):

Yeah , that's what I've done, notice I have added 12 to all three simultaneously, and divided by 2 all together.

OpenStudy (anonymous):

When I put in [3,10] my homework page had a spaz attack

OpenStudy (ash2326):

sorry I made a mistake see this we have -6<=2x-12<=18 add 12 to all sides 6<=2x<=30 divide be 2 the whole inequation 3<=x<=15 As we can see x belongs to [3, 15] do notice that it's closed interval It won't get a attack now:D

OpenStudy (anonymous):

It's fine now haha